OpenStudy (hkmiu):

find (g·f)(3) f(x)= 3x-2 g(x) = x^2

OpenStudy (anonymous):

9

OpenStudy (solomonzelman):

multiply what the functions are equal to together.

OpenStudy (agent0smith):

(g·f)(x) = g(x)·f(x) So you multiply the two functions together: g(x)·f(x) = (3x-2)(x^2) to find (g·f)(3), just plug in x=3 into that^

OpenStudy (hkmiu):

I got 3x^3 - 2x^2 X = 9

OpenStudy (hkmiu):

is that right?

OpenStudy (anonymous):

yes

OpenStudy (hkmiu):

that's funny bc its not one of the answer on my quiz D: ?????????

OpenStudy (anonymous):

what are the answers that you have?

OpenStudy (hkmiu):

http://prntscr.com/2xmra4 ^^^^ screen shot

OpenStudy (hkmiu):

thats not one of them either o.o

OpenStudy (anonymous):

49

OpenStudy (hkmiu):

ohhh wait its a composite

OpenStudy (agent0smith):

3x^3 - 2x^2 plug in x=3, it's not 9.

OpenStudy (hkmiu):

its a composite function ~.~ I read it wrong.

OpenStudy (anonymous):

when you have (g.f)(3) you pluf 3 in f(x) first then the result that you get you plug it into g(x), so answer is 49

OpenStudy (agent0smith):

Oh it's gof(x)... very different to (g·f)(3). g(f(3)) = g(3*3 - 2) = g(7) = 7^2

OpenStudy (hkmiu):

what if you have (fof)(0) and f(x)=x^2-1 wouldn't the outcome still be 0?

OpenStudy (anonymous):

-1

OpenStudy (hkmiu):

ooooooooooo i get it now.

OpenStudy (anonymous):

@agent0smith so was my answer right?

OpenStudy (hkmiu):

so its just f(f(x))

OpenStudy (agent0smith):

^stop posting answers especially when they're wrong @sensuelle1985 f(f(x)) = f(0^2-1) = f(-1) = (-1)^2 - 1 = 0

OpenStudy (hkmiu):

yeah gotta plug -1 into the equation again

OpenStudy (anonymous):

@agent0smith i meant to put 0 its, an error and as far as the other question the answer that I have found was part of her choices!

OpenStudy (hkmiu):

Thank you you guys :3

OpenStudy (anonymous):

and @hkmiu yes you have to put -1 in the equation

OpenStudy (agent0smith):

@sensuelle1985 this wasn't though: You've already chosen the best response. Medals 0 9 25 minutes ago ^ posting answers is not helpful. Posting wrong answers is really unhelpful.

OpenStudy (anonymous):

no i haven't chose any answer as best response! and thats not the point as we are here to help each other out, so cut it Mr smart!!!

OpenStudy (agent0smith):

^don't post answers without showing any working, it's actually part of the site's TOS. Posting an answer, even if correct, doesn't help someone learn.

OpenStudy (anonymous):

thank you for your input! will do for now on